OpenStudy (anonymous):

Will fan and Medal! Solve the equation, x^3x-1=(e^6)^-x

OpenStudy (ooxbrittanyxoo):

\[x=\ln(x)/3(\ln(x)+2)\]

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (ooxbrittanyxoo):

well are you suppose to solve for x?

OpenStudy (anonymous):

I'm assuming...because it only said solve for x.

OpenStudy (ooxbrittanyxoo):

then that is the answer

OpenStudy (anonymous):

I'm sorry i mean that it only said to solve the equation.

OpenStudy (ooxbrittanyxoo):

well im pretty sure that it is still that answer

OpenStudy (anonymous):

The answers possible are, A.) -1/3 B.) 0 C.) 1-9 D.) 7/4

OpenStudy (anonymous):

C is actually 1/9

OpenStudy (trojanpoem):

Can you re-write it using equation button ?

OpenStudy (anonymous):

How do you do it? I don't see the equation button.

OpenStudy (trojanpoem):

\[x^{3x-1}= e^{-6x} \] this ? Down, you will see Equation, draw, Attach file buttons, near "Post" button

OpenStudy (anonymous):

Yes that is it except for the 6 in the power is in () with the e and the -x is still in the power but outside of the ().

OpenStudy (anonymous):

My thing only says, post it doesn't have anything else.

OpenStudy (trojanpoem):

(2^2)^2 = 2^4 so the powers are multiplied.

OpenStudy (anonymous):

It's like e^6 are in ( ) and then -x is ^the power but outside of the ( )